A newspaper is known to have a mean number of 4 typing errors per page. Find the probability that;
i. at most four typographical errors will be found on a page
ii. more than two typographical errors will be found on a page


Answer by Boreal(15235)

Poisson distribution--random, could theoretically be infinite, count data, independent occurrences.
More than 2 is 1- p(2)-p(1)-p(0)
p(0)=e^(-4)*4^0/0!=e^(-4)=0.0183
p(1)=e^(-4)*4^1/1!=0.0733
p(2)=e^(-4)*4^2/2!=0.1465
p(3)=e^(-4)*64/6=0.1954
p(4)=e^(-4)*256/24=0.1954
At most 4 is the sum of these 5, or 0.6289
More than two is 1-0.0183-0.0733-0.1465=0.7619